OpenStudy (anonymous):

solve for x 2 ln3x=5

OpenStudy (anonymous):

ln 3x = 5/2 3x = e^(5/2) x = e^(5/2) / 3

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

i could have sworn i did this problem not 3 minutes ago...

OpenStudy (anonymous):

ha ha ha, ITS ALL A BLUR :D

OpenStudy (anonymous):

\[x=\frac{e^\frac{5}{2}}{3}\], yes...